OpenStudy (gabylovesyou):

What is the arc length of arc CD in the circle below? Circle A is shown with a radius labeled 6 feet and a central angle marked 30 degrees. 3.14 feet 4.88 feet 6.98 feet 11.78 feet

OpenStudy (gabylovesyou):

@campbell_st

OpenStudy (campbell_st):

find the cricumference of the circle \[C = 2\times \pi \times r\]

OpenStudy (gabylovesyou):

25.68

OpenStudy (campbell_st):

then the fraction of the circumference is \[\frac{30}{360}\]. Just multiply the circumference by the fraction

OpenStudy (gabylovesyou):

2.14

OpenStudy (campbell_st):

ummm I think you have a typo...

OpenStudy (gabylovesyou):

circumfrence is 2 * 2.14 * 6 = 25.68 25.60 * (30/360) = 2.14

OpenStudy (gabylovesyou):

thats what i get.. ;/

OpenStudy (campbell_st):

ummm I think you'll find that pi = 3.14

OpenStudy (gabylovesyou):

ohhhhh my bad cx rofl

OpenStudy (gabylovesyou):

3.14 .. is the answer

OpenStudy (gabylovesyou):

thank you

OpenStudy (campbell_st):

correct